solve using the addition method.
2x - 3y = -6
x + y = 7

Answer by stanbon(75887)

solve using the addition method.
2x - 3y = -6
x + y = 7
------------
multiply thru the 2nd eqution by 2 to get
2x+2y = 14
-------
Subtract this from the 1st equation to get:
-5y = -20
y = 4
-------------------
Substitute that into x+y = 7
x + 4 = 7
x = 3
------------
Solution :(3,4)
==================
Cheers,
Stan H.